1. Find the points of horizontal and vertical tangency (if any) to the polar curve r = 2 – 2 cos(0)
Answer #1

r = 2-2 Cane x=rCaro = (2-2 Caro) Gaxo 2 Ceen o - 2 Cax 20 y = -r Sino = (2-2 600) Sino 2 = 2 Sino 2 Sino Gno da = - 2 Sino -at o, o dx do - Sin 0 + 4 Smo Cono dx do o, is not a solution at 0, - 211 3 da do 2 sm 21 3 +4 Som 2 11 Coming dx do - 253 0at Og = 0 - 26x²o - 26x²0 + 2 sm²0 dy do dy do =o o il net a solution at Os - TT dy = 2 lant - 2 Cax²n + 2 fm² do dy --4 → ©